# Remote electric-field control of antiferromagnetic magnon-polaritons

### v1

@@ 3.2 Liquid Crystal Cell
We present a schematic of the liquid crystal cell cross-section in Fig. 5(b). We use the liquid crystal cell functionalized with a highly birefringent nematic liquid crystal mixture 1825, which is characterized by ordinary and extraordinary refractive indices of and , respectively41, 42. We realize a planar cell arrangement, where a 100 m thick layer of liquid crystal is sandwiched between two pristine slabs of fused quartz (Silica), each 1500 m thick. The surfaces of the silica slabs facing the liquid crystal are coated with a thin film of a conducting polymer Electra 92 (AR-PC 5090.02 by ) and uniformly rubbed with a 100 nm-thick PMMA resist. The PMMA film served as an alignment layer, facilitating the alignment of the liquid crystal molecules near its surface in the direction of rubbing. We achieve planar alignment of the liquid crystal in its bulk by matching the rubbing directions at the opposite sides of the cell.
@@
The coating parameters for each layer are as follows: first, the silica glasses are covered with the Electra 92 via spin-coating (2000 rpm for 60 s) and baked on a hotplate at 105∘C for 5 minutes. According to the product information, the resulting coated film should have a thickness of 60 nm. Next, a 100 nm layer of PMMA resist (200K, 2 by ) is deposited on top of the Electra 92 layer by spin-coating (6000 rpm for 45 s). The resist is baked on a hotplate at 150∘C for 10 minutes. Electra is a conductive coating used in e-beam lithography, and it is semi-transparent in the THz range, so this layered structure (Electra-PMMA) forms a THz semi-transparent electrode.
@@
3.3 Numerical calculations
We perform qualitative modeling of electromagnetic field distributions of the Fabry-Pérot cavity modes using the transfer matrix method. We calculate the reflection from a system composed of parallel-plane slabs of silica glass, liquid crystal, NiO, and a gap filled with air between NiO and the liquid crystal cell. We neglect the thin layers of PMMA and Electra 92 in our calculations, as well as reflections from the copper mirror. We normalize the -field for 30 V for each mode to the -field at 0 V in silica glass. We assume thicknesses and refractive indices: silica 1.5 mm , liquid crystal 0.1 mm , and NiO 0.51 mm . In this analysis, we focus on modes that remain unaffected by interactions with the magnon mode, setting = 1.
4 Acknowledgements
The work was supported by the European Union through the ERC-ADVANCED grant TERAPLASM (No. 101053716). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge the support of the "Center for Terahertz Research and Applications (CENTERA2)" project (FENG.02.01-IP.05-T004/23) carried out within the "International Research Agendas" program of the Foundation for Polish Science co-financed by the European Union under the European Funds for a Smart Economy Programme. This work was partially supported by Pasific2 of the Polish Academy of Sciences, sponsored by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No.847639 and from the Ministry of Education and Science of Poland. This work was also partially supported by the Sonata bis-13 2023/50/E/ST3/00584 grant of the National Science Centre of Poland. R. K. and O. S. acknowledge the financial support from the National Science Centre under grant SONATA BIS-12 No. 2022/46/E/ST7/00454.

### v2

@@ 4.3 Liquid Crystal Cavity
We present a schematic of the liquid crystal cell cross-section in Fig. 1(d). We use the liquid crystal cell functionalized with a highly birefringent nematic liquid crystal mixture 1825, which is characterized by ordinary and extraordinary refractive indices of and , respectivelyRKowerdziej_APL2013, MReuter_APLM2013. We realize a planar cell arrangement, where a 100 m thick layer of liquid crystal is sandwiched between two pristine slabs of fused quartz (Silica), each 1500 m thick. The surfaces of the silica slabs facing the liquid crystal are coated with a thin film of a conducting polymer, Electra 92 (AR-PC 5090.02 by ), and uniformly rubbed with a 100 nm-thick PMMA resist. The PMMA film served as an alignment layer, facilitating the alignment of the liquid crystal molecules near its surface in the direction of rubbing. We achieve planar alignment of the liquid crystal in its bulk by matching the rubbing directions at the opposite sides of the cell.
@@
The coating parameters for each layer are as follows: first, silica glasses are covered with Electra 92 via spin-coating (2000 rpm for 60 s) and then baked on a hotplate at 105∘C for 5 minutes. According to the product information, the resulting coated film should have a thickness of 60 nm. Next, a 100 nm layer of PMMA resist (200K, 2 by ) is deposited on top of the Electra 92 layer by spin-coating (6000 rpm for 45 s). The resist is baked on a hotplate at 150∘C for 10 minutes. Electra is a conductive coating used in e-beam lithography, and it is semi-transparent in the THz range, so this layered structure (Electra-PMMA) forms a THz semi-transparent electrode.
@@
4.4 Simulations
We perform qualitative simulations of electromagnetic field distributions of the Fabry-Perot cavity modes using the transfer matrix method. We calculate the reflection from a system composed of parallel-plane slabs of silica glass, liquid crystal, NiO (without magnon resonance), and a gap filled with air between NiO and the liquid crystal cell. We neglect the thin layers of PMMA and Electra 92 in our calculations. We normalize the -field for 30 V for each mode to the -field at 0 V in silica glass. We assume thicknesses and refractive indices: silica 1.5 mm , liquid crystal 0.1 mm in the range of to , gap between the liquid crystal cell and NiO of 0.91 mm, and NiO thickness of 335 m and refractive index . In this analysis, we focus on modes that remain unaffected by interactions with the magnon mode, setting = 1. We use a thermocouductive paste to fix the NiO on a copper plate, which is necessary for precise temperature control. We experimentally find that the reflection spectra from NiO placed on metal, with and without paste, are very distinct. This is because the paste we use has strong absorption and relatively small reflectivity in the THz range, which we approximate in our model as a 20 m-thick layer with .
In Fig. 6, we simulate reflection spectra with the magnon. Based on the NiO transmission result, we use the Lorentzian to describe the NiO magnon MBialek_AFM2024, ie
where is temperature-dependent magnon frequency, is its oscillator strength, and GHz is its width. We sweep in the range of 0.91 to 0.98 THz. Detailed frequency shifts of modes by about 10 GHz are most likely caused by approximations in the thicknesses of certain layers and their refractive indices.
Associated Content
The data that support the findings of this study are available in RepOD at https://doi.org/10.18150/WPQCOV
Acknowledgements
The authors thank Prof. W. Pacuski, and Prof. J.-Ph. Ansermet for valuable discussions.
Funding Sources
The European Union supported the work through the ERC-ADVANCED grant TERAPLASM (No. 101053716). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge the support of the "Center for Terahertz Research and Applications (CENTERA2)" project (FENG.02.01-IP.05-T004/23) carried out within the "International Research Agendas" program of the Foundation for Polish Science, co-financed by the European Union under the European Funds for a Smart Economy Programme. This work was partially supported by Pasific2 of the Polish Academy of Sciences, sponsored by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847639 and by the Ministry of Education and Science of Poland. M. B. acknowledges the financial support of the Sonata BIS-13 No. 2023/50/E/ST3/00584 grant of the National Science Centre of Poland. R. K. and O. S. acknowledge the financial support from the National Science Centre under grant SONATA BIS-12 No. 2022/46/E/ST7/00454.
Supporting Information
Supporting Information is available: S1. Time-domain reflection spectra; S2. NiO magnon resonance; S3. Spectra deconvolution; S4. Voltage bias effect on cavity modes.